OpenStudy (anonymous):

What is the solution to the equation 3^(–2x) ≈ 4 ?

OpenStudy (anonymous):

options: x = –0.631 x = –0.396 x = 0.396 x = 0.631

OpenStudy (anonymous):

take the log of both \[ln(3^{-2x}=ln(4)\]

OpenStudy (anonymous):

you can now pull out the -2x from the top \[-2xln(3)=ln(4)\]

OpenStudy (anonymous):

divide by -2ln(3

OpenStudy (anonymous):

okay , then what do i do ?

OpenStudy (anonymous):

\[x=\frac{ln(4)}{-2ln(3)}\]

OpenStudy (anonymous):

okay , im not really sure what the heck your doing nd idk how to move on .

OpenStudy (anonymous):

\[ln(a^b)=bln(a)\]

OpenStudy (anonymous):

omgggg ! what are you doinggg ?

OpenStudy (anonymous):

i'm using the property \[ln(a^b)=bln(a)\]

OpenStudy (anonymous):

okayy , whats the answerrrr ? :P

OpenStudy (anonymous):

nvm , got it !

OpenStudy (anonymous):

alright. Sorry about not responding. I was busy with something lse